OpenStudy (anonymous):

please help foh h4p2 what is The minimum value of RL in Ohms, that guarantees that the circuit will operate this way?

OpenStudy (anonymous):

mine was 524.934ohms

OpenStudy (anonymous):

it maybe different for u depeniding on the values

OpenStudy (anonymous):

What are the givven values for the task?

OpenStudy (anonymous):

guys for VI=11.0V, RIN=1 kΩ, and vi=40.0mV. min RL should be what ?

OpenStudy (anonymous):

@moyani .. for the ckt without zener ur values are for RL=2K Vo= 22/3 vo= 2/75 for RL=4K Vo= 44/5 vo= 4/125

OpenStudy (anonymous):

@moyani ... are they correct ?? pls lemme know !!

OpenStudy (anonymous):

Try 850 for minimum RL and tell is it ok?

OpenStudy (anonymous):

yah guys you're the best... Correct